Americans drive about 4.0 x 10^9 miles per day, and their cars get an average of 20 miles per gallon of gas used. For each kilogram of gasoline used, about 3 kg of carbon dioxide is produced. How many kilograms of CO2 are emitted into the atmosphere by cars in the US each day? [1 gal = 3.5 kg]
Answer #1

No of gallons of gasoline used =4x10^9/20

=2x10^8

=2x10^8x3.5 Kg

=7x10^8 Kg

1 Kg gasoline produces 3 Kg CO2

So 7x10^8 Kg gasoline produces (7x10^8)(3)

=21x10^8 Kg CO2

=2.1x10^9 Kg
Answer #2

20 miles covered per gallon of gas

So, for 4.0 x 10^9 miles, 4/20 x 10^9 gallon of gas = 0.2  x 10^9 gallon of gas

1 gallon= 3.5kg

So, 0.2  x 10^9 gallon = 3.5 x  0.2  x 10^9 kg = 0.7 x 10^9 kg gas

1kg of gasoline, 3kg CO2

So for  0.7 x 10^9 kg gas, we get  0.7 x 10^9  x 3kg CO2 = 2.1x 10^9 kg gas

each kilogram of gasoline used, about 3 kg of carbon dioxide is produced
Answer #3

total gallons of gas = 4.0 X 10^9 / 20 = 2.0 X 10^8 gallons

mass of gas = 2.0 X 10^8 X 3.5 = 7.0 X 10^8 kg

so mass of CO2 = 7.0 X 10^8 X 3 = 2.1 X 10^9 kg